OpenStudy (anonymous):

some please work this problem completely step by step show ur work so i can follow along divide & simplify z^2-49 divided by 2z-14 ------- --------- z^2-8z+16 z^2-2z-63 i have done all the way till this part where i am stuck (z−9)(z+7)^2 ------------ 2(z−4)^2

OpenStudy (anonymous):

\[\frac{z^2-49}{2z-14}\] \[\frac{(z-7)(z+7)}{2(z-7)}\] \[\frac{z+7}{2}\]

OpenStudy (anonymous):

\[\frac{z^2-49}{z^2-8z+16}\div \frac{2z-14}{z^2-2z-63}\] multiply by reciprocal \[\frac{z^2-49}{z^2-8z+16}\times \frac{z^2-2z-63}{2z-14}\] factor \[\frac{(z+7)(z-7)}{(z-4)(z-4)}\times \frac{(z-9)(z+7)}{2(z-7)}\] cancel the factors that will cancel \[\frac{(z+7)^2(z-9)}{2(z-4)^2}\] that is your final answer. It can't be reduced or simplified anymore.